OpenStudy (anonymous):

How do I find the range of the function f(x)=tan(x)-1 if the domain is [-π/2, π/2] ?

OpenStudy (anonymous):

Domain: numbers you feed the function. Range: numbers you get out. Lets try \[f(\pi/2)=\tan(\pi/2)-1\], eek! that is \[\frac{\sin\pi/2}{\cos\pi/2}-1\] which is dividing by ZERO! ... better use the limit from the left to stop the world from imploding.

OpenStudy (anonymous):

repeat for \[-\pi/2\] and you've got the range of f(x).